OpenStudy (anonymous):

Show that waves shone directly upward, with k-vector = k z-hat, will reflect back provided w < wp max.

OpenStudy (anonymous):

\[w p = \sqrt{e^2 \frac{ N(z) }{ \epsilon0*m }}\], where e = charge of an electron in C, m = mass of an electron in Kg, eps0 = permitivvity of free space, N(z) = particle density

OpenStudy (anonymous):

\[N(z,w) = \sqrt{1 - w p^2(z)/w^2}\]

OpenStudy (anonymous):

So the first N(z) = N(z,w)